A die means a standard cube whose faces are numbered 1, ..., 6 and each has equal probability of being rolled. A green die and a red die are rolled simultaneously, with equal probability for each of the 36 possible outcomes. Describe the sample space. Then find the probability that: a. the sum is 7;

b the sum is 11;

thesumis7or11;

thesumisneither7nor11;

b. the same value occurs on each die;

c. the value on the green die is greater than 5;

d. the value on each die is less than 6;

e. the sum is even.
Answer #1
red/green 1 2 3 4 5 6
1 11 21 31 41 51 61
2 12 22 32 42 52 62
3 13 23 33 43 53 63
4 14 24 34 44 54 64
5 15 25 35 45 55 65
6 16 26 36 46 56 66

total possible outcomes =36

a) sum is 7
possible outcomes are(1,6)(2,5)(3,4)(4,3)(5,2)(6,1)

so,P(sum is 7)=6/36 = 1/6

b) sum is 11
possible outcomes are(5,6)(6,5)

so ,P(sum is 11) = 2/36

c) P(sum is 7 or 11)=P(sum is 7)+P(sum is 11) = 6/36+2/36 = 8/36 = 2/9

d)P( sum is neither 7 or nor 11 ) = 1-P(sum is 7 or 11) = 1-8/36 = 28/36 = 7/9
____________________________________________________________________________________

b. the same value occurs on each die;

possible outcomes are (11)(22)(33)(44)(55)(66)

so, probability = 6/36 = 1/6

c) value on green die will be 6 and red die can take any value
possible outcomes are = (1,6)(2,6)(3,6)(4,6)(5,6)(6,6)
P(value on green die is greater than 5)=6/36 = 1/36

d)

possible outcomes are

11 21 31 41 51
12 22 32 42 52
13 23 33 43 53
14 24 34 44 54
15 25 35 45 55

so, P(value on each die is less than 6) = 25/36

e)

possible outcomes are

11 31 51
22 42 62
13 33 53
24 44 64
15 35 55
26 46 66

P(sum is even) = 18/36 = 1/2